### v1

# An elementary proof of a universal approximation theorem

Abstract
In this short note, we give an elementary proof of a universal approximation theorem for neural networks with three hidden layers and increasing, continuous, bounded activation function. The result is weaker than the best known results, but the proof is elementary in the sense that no machinery beyond undergraduate analysis is used.
1 Introduction
There are several versions of universal approximation theorems known, including the very well-known ones from [1, 2, 3]. Each of them states that some collection of neural networks is dense in some space of continuous functions with respect to the uniform norm. In this short note, we present what we believe to be a new and atypically elementary proof of one such theorem. If $\sigma$ is a 0-1 squashing function (a.k.a. a sigmoidal function), we show that the collection of neural networks with three hidden layers and activation function $\sigma$ (except at the output) is dense in the space $C(K)$ of real-valued continuous functions on a compact set $K\subset\mathbb{R}^{n}$. The result given here is weaker than the best known results, but the argument relies only on basic results about compact sets and continuous functions which are generally covered in an undergraduate analysis course; it is really nothing more than an exercise in “epsilon chasing,” though the underlying intuitive motivation is fairly natural.
2 Notations
For the entirety of this note, we assume the following:
  1. 1.
    $\sigma:\mathbb{R}\longrightarrow\mathbb{R}$ is increasing, continuous, and $\lim_{x\to\-\infty}\sigma(x)=0$ and $\lim_{x\to\infty}\sigma(x)=1$.
  2. 2.
    $K$ is a compact subset of $\mathbb{R}^{n}$.
For example, the sigmoid function $\sigma(x)=1/(1+e^{-x})$ satisfies the first assumption. A function satisfying that assumption will be called a 0-1 squashing function. There’s nothing particularly special about 0 and 1 here. The crucial properties used in our arguments are really just that $\sigma$ is increasing, continuous, bounded, and non-constant. But the discussion is simplified with these additional limiting assumptions.
The compact set $K$ will represent the domain of a function to be approximated. In particular, the unit cube $K=[0,1]^{n}$ is a special case.
Of course, it’s not at all necessary for a neural network to use the same activation function in each hidden layer, but that will suffice for our purposes. Fix a positive integer $n$ and let
$\mathcal{N}_{1}$ $=$ $\{f\in C(K)\,\,:\,\,f(x_{1},\ldots,x_{n})=a_{0}+a_{1}x_{1}+\cdots+a_{n}x_{n},$
[…]
$\mathcal{N}_{1}^{\sigma}$ $=$ $\{F\in C(K)\,\,:\,\,F=\sigma\circ f,\hskip 12.0pt\mbox{ for some }f\in\mathcal{N}_{1}\}.$
That is, $\mathcal{N}_{1}$ is the set of all affine functions of $x_{1},\ldots,x_{n}$. And $\mathcal{N}_{1}^{\sigma}$ is the set of all possible node output functions in Layer 1. For $k\geq 1$, we define
$\mathcal{N}_{k+1}$ $=$ $\{g\in C(K)\,\,:\,\,g=a_{0}+a_{1}F_{1}+\cdots+a_{m}F_{m},$
[…]
$\mathcal{N}_{k+1}^{\sigma}$ $=$ $\{G\in C(K)\,\,:\,\,G=\sigma\circ g,\hskip 12.0pt\mbox{ for some }g\in\mathcal{N}_{k+1}\}.$
Thus, $\mathcal{N}_{k+1}^{\sigma}$ is the set of all possible node output functions in Layer $k+1$. It follows easily from the definitions above that $\mathcal{N}_{k}^{\sigma}\subset\mathcal{N}_{k+1}$, and if $g_{1},g_{2}\in\mathcal{N}_{k}$ then $(a_{0}+a_{1}g_{1}+a_{2}g_{2})\in\mathcal{N}_{k}$ for all $a_{0},a_{1},a_{2}\in\mathbb{R}$. We will use these facts frequently in the sequel.
3 Three separation lemmas
Roughly speaking, the strategy of the proof is to first show that $\sigma$ separates points in $\mathbb{R}^{n}$ in a strong sense. We then show that $\mathcal{N}_{2}$ separates points from closed sets in a similarly strong sense, and that $\mathcal{N}_{3}$ separates closed sets from closed sets again in a strong sense. Finally, the theorem will be proven using this result about $\mathcal{N}_{3}$.
Lemma 3.1
Let $x_{0}$ and $x_{1}$ be distinct real numbers. For each $\epsilon\in(0,1)$ there exist $s,t\in\mathbb{R}$ such that $\sigma(s+tx_{0})<\epsilon$ and $\sigma(s+tx_{1})>1-\epsilon$.
  • Proof:
    By definition of a 0-1 squashing function, there exist $y_{0},y_{1}\in\mathbb{R}$ such that $\sigma(y_{0})=\epsilon/2$ and $\sigma(y_{1})=1-\epsilon/2$. Since $x_{1}-x_{0}\neq 0$, the linear system
[…]
    has a solution, say $(s_{0},t_{0})^{T}$. Thus,
    $\sigma(s_{0}+t_{0}x_{0})$ $=$ $\sigma(y_{0})=\epsilon/2<\epsilon,\hskip 12.0pt\mbox{ and }$
    $\sigma(s_{0}+t_{0}x_{1})$ $=$ $\sigma(y_{1})=1-\epsilon/2>1-\epsilon.$
The next lemma asserts that we can separate points from closed sets using functions in $\mathcal{N}_{2}$ in a strong sense. That is, one hidden layer suffices to separate points from closed sets in the described sense.
Lemma 3.2
Let $B\subset K$ be a closed set, and $\boldsymbol{x}_{0}\in K-B$. For each $\epsilon>0$ there exists $g\in\mathcal{N}_{2}$ such that $g>1-\epsilon$ on $B$ and $g(\boldsymbol{x}_{0})<\epsilon$.
  • Proof:
    Without loss of generality, assume $30<\epsilon<1/$ 3. Let ${\boldsymbol{b}}\in B$. Since ${\boldsymbol{b}}\neq\boldsymbol{x}_{0}$, there exists $f_{\boldsymbol{b}}\in\mathcal{N}_{1}$ such that $f_{\boldsymbol{b}}(\boldsymbol{x}_{0})<\epsilon/2$ and $f_{\boldsymbol{b}}({\boldsymbol{b}})>1-\epsilon/2$. Let
    $U_{\boldsymbol{b}}=\{\boldsymbol{x}\in K\,\,:\,\,f_{\boldsymbol{b}}(\boldsymbol{x})>1-\epsilon\}.$
    Since $f_{\boldsymbol{b}}$ is continuous, $U_{\boldsymbol{b}}$ is open and ${\boldsymbol{b}}\in U_{\boldsymbol{b}}$, so $\{U_{\boldsymbol{b}}\}_{{\boldsymbol{b}}\in B}$ is an open cover of the compact set $B$. Let ${\boldsymbol{b}}_{1},\ldots,{\boldsymbol{b}}_{N}\in B$ such that $\{U_{{\boldsymbol{b}}_{1}},\ldots,U_{{\boldsymbol{b}}_{N}}\}$ is a cover of $B$.
    By Lemma 3.1 and using the fact that $\sigma$ is monotonic, there exist $s,t\in\mathbb{R}$ such that $\sigma(s+tx)<\epsilon/N$ on $(-\infty,\epsilon)$ and $\sigma(s+tx)>1-\epsilon$ on $(1-\epsilon,\infty)$. For each $1\leq j\leq N$, let $F_{j}=\sigma(s+tf_{{\boldsymbol{b}}_{j}})$. Then $F_{j}\in\mathcal{N}_{1}^{\sigma}$ and $F_{j}(\boldsymbol{x}_{0})<\epsilon/N$ and $F_{j}\geq 1-\epsilon$ on $U_{{\boldsymbol{b}}_{j}}$.
    Let $g=\sum_{j=1}^{N}F_{j}$, and it follows that $g\in\mathcal{N}_{2}$, $g(\boldsymbol{x}_{0})<\epsilon$ and $g\geq 1-\epsilon$ on $B$.
The following lemma is the major tool in the proof of the universal approximation theorem we will present. It asserts that two hidden layers suffice to separate disjoint closed sets, in a sense very similar to the previous lemma. Its proof is quite similar to the previous one.
Lemma 3.3
Let $A$ and $B$ be disjoint closed subsets of $K$. For each $\epsilon>0$ there exists $h\in\mathcal{N}_{3}$ such that $h>1-\epsilon$ on $A$ and $h<\epsilon$ on $B$.
  • Proof:
    Again without loss of generality we assume $\epsilon\in(0,1/3)$. For each ${\boldsymbol{a}}\in A$, by Lemma 3.2, there exists $\widetilde{g}_{\boldsymbol{a}}\in\mathcal{N}_{2}$ such that $\widetilde{g}_{\boldsymbol{a}}>1-\epsilon/2$ on $B$ and $\widetilde{g}_{\boldsymbol{a}}({\boldsymbol{a}})<\epsilon/2$. Let $g_{\boldsymbol{a}}=1-\widetilde{g}_{\boldsymbol{a}}$, so that $g_{\boldsymbol{a}}\in\mathcal{N}_{2}$, and $g_{\boldsymbol{a}}<\epsilon/2$ on $B$, and $g_{\boldsymbol{a}}({\boldsymbol{a}})>1-\epsilon/2$. Let
    $U_{\boldsymbol{a}}=\{\boldsymbol{x}\in K\,\,:\,\,g_{\boldsymbol{a}}(\boldsymbol{x})>1-\epsilon\}.$
    Since $g_{\boldsymbol{a}}$ is continuous, each $U_{\boldsymbol{a}}$ is open. And since ${\boldsymbol{a}}\in U_{\boldsymbol{a}}$, it follows that $\{U_{\boldsymbol{a}}\}_{{\boldsymbol{a}}\in A}$ is an open cover of the compact set $A$. Let ${\boldsymbol{a}}_{1},\ldots,{\boldsymbol{a}}_{N}\in A$ for which $\{U_{{\boldsymbol{a}}_{1}},\ldots,U_{{\boldsymbol{a}}_{N}}\}$ is a cover of $A$.
    Let $s,t\in\mathbb{R}$ for which $\sigma(s+tx)<\epsilon/N$ on $(-\infty,\epsilon)$ and $\sigma(s+tx)>1-\epsilon$ on $(1-\epsilon,\infty)$. Set
    $h=\sum_{j=1}^{N}\sigma(s+tg_{{\boldsymbol{a}}_{j}}).$
    First note that $h\in\mathcal{N}_{3}$. If ${\boldsymbol{a}}\in A$, then ${\boldsymbol{a}}\in U_{{\boldsymbol{a}}_{k}}$ for some $k$, which implies that $g_{{\boldsymbol{a}}_{k}}({\boldsymbol{a}})>1-\epsilon$ and so $h({\boldsymbol{a}})>1-\epsilon$. And if ${\boldsymbol{b}}\in B$ it follows that $g_{{\boldsymbol{a}}_{j}}({\boldsymbol{b}})<\epsilon/2$ for all $j$, which gives $\sigma(s+g_{{\boldsymbol{a}}_{j}}({\boldsymbol{b}}))<\epsilon/N$ for all $j$ and hence $h({\boldsymbol{b}})<\epsilon$.
Recall that if $K$ is a compact subset of $\mathbb{R}^{n}$ and $f:K\longrightarrow\mathbb{R}$ is continuous, then the sup norm of $f$, denoted $\|f\|_{u}$, is
$\|f\|_{u}=\sup\{|f(\boldsymbol{x})|\,\,:\,\,\boldsymbol{x}\in K\}.$
Theorem 3.4 (A Universal Approximation Theorem)
Let $\sigma$ be a 0-1 squashing function, and $\mathcal{N}_{k},\mathcal{N}_{k}^{\sigma}$ as previously defined. Let $T:K\longrightarrow\mathbb{R}$ be a continuous function. For each $\epsilon>0$ there exists $f\in\mathcal{N}_{4}$ such that $\|f-T\|_{u}<\epsilon$; that is, $\mathcal{N}_{4}$ is dense in $C(K)$ with respect to the $\sup$ norm.
  • Proof:
    By way of contradiction, suppose that $T:K\longrightarrow\mathbb{R}$ is a continuous function and
    $\inf_{f\in\mathcal{N}_{4}}\|f-T\|_{u}=\alpha>0.$
    Let $\widehat{f}\in\mathcal{N}_{4}$ with $\alpha\leq\|\widehat{f}-T\|_{u}<4\alpha/3$. Define
    $U^{+}$ $=$ $\left\{\boldsymbol{x}\in K\,\,:\,\,\frac{\alpha}{3}\leq(\widehat{f}-T)(\boldsymbol{x})\leq\frac{4\alpha}{3}\right\},$
    $U^{-}$ $=$ $\left\{\boldsymbol{x}\in K\,\,:\,\,-\frac{4\alpha}{3}\leq(\widehat{f}-T)(\boldsymbol{x})\leq-\frac{\alpha}{3}\right\}.$
    Since $U^{+},U^{-}$ are closed and disjoint, by Lemma 3.3 there exists $H\in\mathcal{N}_{3}^{\sigma}$ such that
    $0\leq H<\frac{1}{6}\hskip 12.0pt\mbox{ on }U^{-},\hskip 12.0pt\mbox{ and }\frac{5}{6}<H\leq 1\hskip 12.0pt\mbox{ on }U^{+}.$
    Consider $f=\widehat{f}-\alpha H+\frac{\alpha}{2}\in\mathcal{N}_{4}$. We claim that $\|f-T\|_{u}<\alpha$.
    Suppose first that $\boldsymbol{x}\in U^{+}$. Then
    $(f-T)(\boldsymbol{x})$ $=$ $(\widehat{f}-T)(\boldsymbol{x})-\alpha H(\boldsymbol{x})+\frac{\alpha}{2}$
    $<$ $\frac{4\alpha}{3}-\frac{5\alpha}{6}+\frac{\alpha}{2}=\alpha.$
    We also have that
    $(f-T)(\boldsymbol{x})$ $=$ $(\widehat{f}-T)(\boldsymbol{x})-\alpha H(\boldsymbol{x})+\frac{\alpha}{2}$
    $\geq$ $\frac{\alpha}{3}-\alpha+\frac{\alpha}{2}=-\frac{\alpha}{6}>-\alpha,$
    And therefore $|f-T|<\alpha$ on $U^{+}$. Suppose now that $\boldsymbol{x}\in U^{-}$. Then
    $(f-T)(\boldsymbol{x})$ $=$ $(\widehat{f}-T)(\boldsymbol{x})-\alpha H(\boldsymbol{x})+\frac{\alpha}{2}$
    $\leq$ $-\frac{\alpha}{3}+\frac{\alpha}{2}=\frac{\alpha}{6}<\alpha,$
    and
    $(f-T)(\boldsymbol{x})$ $=$ $(\widehat{f}-T)(\boldsymbol{x})-\alpha H(\boldsymbol{x})+\frac{\alpha}{2}$
    $>$ $-\frac{4\alpha}{3}-\frac{\alpha}{6}+\frac{\alpha}{2}=-\alpha,$
    and hence $|f-T|<\alpha$ on $U^{-}$. Finally, suppose that $\boldsymbol{x}\in K-(U^{-}\cup U^{+})$. Since $-\frac{4\alpha}{3}<(\widehat{f}-T)(\boldsymbol{y})<\frac{4\alpha}{3}$ for all $\boldsymbol{y}\in K$, it follows that $-\frac{\alpha}{3}<(\widehat{f}-T)(\boldsymbol{x})<\frac{\alpha}{3}$. From this and $0\leq H\leq 1$ it follows that
    $(f-T)(\boldsymbol{x})$ $=$ $(\widehat{f}-T)(\boldsymbol{x})-\alpha H(\boldsymbol{x})+\frac{\alpha}{2}$
    $<$ $\frac{\alpha}{3}+\frac{\alpha}{2}=\frac{5\alpha}{6}<\alpha,$
    and
    $(f-T)(\boldsymbol{x})$ $=$ $(\widehat{f}-T)(\boldsymbol{x})-\alpha H(\boldsymbol{x})+\frac{\alpha}{2}$
    $>$ $-\frac{\alpha}{3}-\alpha+\frac{\alpha}{2}=-\frac{5\alpha}{6}>-\alpha.$
    But $f\in\mathcal{N}_{4}$ and $\|f-T\|_{u}<\alpha$ is a contradiction, and so the theorem is proven.
3.1 Some extensions of Theorem 3.4
There are a number of easily deduced corollaries that extend the result. For example, if $\sigma$ is strictly increasing, we can obtain a similar result using $\mathcal{N}_{4}^{\sigma}$ for continuous functions from $K$ to the image $\sigma(\mathbb{R})$. Of course another corollary is that functions $C(K,\mathbb{R}^{m})$ can be uniformly approximated by functions in $(\mathcal{N}_{4})^{m}$, and these two extensions could also be combined. But these are all very standard and obvious extensions which apply quite generally.